OpenStudy (anonymous):

Please solve giving step-by-step detail: lim x-> pi/2 (tanx(^cosx))

OpenStudy (anonymous):

Let A=lim x->pi/2 (tanx)^cosx. Put both sides inside ln function: \[\ln A=\ln[\lim_{x \rightarrow \frac{ \pi }{ 2 }}\tan x ^{\cos x}]\] Limit and ln function can be replaced: \[\ln [\lim_{x \rightarrow \frac{ \pi }{ 2 }} \tan x ^{\cos x}]=\lim_{x \rightarrow \frac{ \pi }{ 2 }}[\ln (\tan x)^{\cos x}]\] Use now property of ln function: \[\ln Y ^{Z}=Z*\ln Y\] \[\lim_{x \rightarrow \frac{ \pi }{ 2 }}[(\cos x)*\ln \tan x]\]. You can write this as: \[\lim_{x \rightarrow \frac{ \pi }{ 2 }}[\frac{ \ln \tan x }{ \cos x}]\] These form is inf/inf and you can aply l'hopital rule: \[=\lim_{x \rightarrow \frac{ \pi }{ 2 }}\frac{ \frac{ 1 }{ \tan x* \cos ^{2} x} }{ \frac{ \sin x }{ \cos ^{2}x }}\] Simplyfing these:\[\lim_{x \rightarrow \frac{ \pi }{ 2 }}\frac{ \cos x }{ \sin ^{2} x}=0\] And now recall that ln(A)=0, where A is solution to given limit. Function ln(Y) is equal to 0 only if its argument Y is equal to 1, so we get that A=1 and solution to given limit is 1.

OpenStudy (anonymous):

How did u get (ln tanx)/cos x ??

OpenStudy (anonymous):

I made mistake in typing, correct is:\[\frac{ \ln \tan x }{ \frac{ 1 }{cosx} }\]